a manufacturer wants to build a spring that takes a force 10 n(in negative direction) to compress it 0.2 m from the equilibrium position.the spring should be able to stretch 0.5 m from the equilibrium position.

as a mechanical engineer, you were asked to present how much work should be done to stretch the spring.the presentationshall include important components of the problems, complete and correct computations and a logical and organizedexplanation.


1
Expert's answer
2021-04-29T17:00:02-0400

"F=-10N, x_1=0.2m,\n\n\n\nx_2=0.5m"


Stretch in spring "x=x_1-x_2=0.2-0.5=-0.3m"


Work done to stretch the string "W=Fx=-10\times -0.3=3J"